# Robots that Evolve with Us: Modular Co-Design for Personalization, Adaptability, and Sustainability

### v1

@@ 1. Introduction
Robots are increasingly envisioned not as short-term tools but as long-term companions that support people throughout their lives (Leite et al., 2013; Kaplan, 2005; Chen et al., 2025a). From childhood education to workplace productivity and eldercare (Chen et al., 2025a; Bemelmans et al., 2012; Kory-Westlund and Breazeal, 2019; Westlund et al., 2018), socially assistive robots hold the promise of providing functional help (Deutsch et al., 2019; Klamer and Allouch, 2010), emotional connection (Šabanović et al., 2013), and continuity of care across the human lifespan (Irfan et al., 2023; Kidd and Breazeal, 2008). Realizing this promise, however, requires moving beyond one-size-fits-all systems. Because people’s needs, values, and contexts shift over time (Carstensen et al., 1999; Fung et al., 2001), robots should remain not only useful at one life stage but also adaptable over decades (De Graaf et al., 2015). This raises a central challenge for Human–Robot Interaction (HRI): how can robots remain relevant, meaningful, and sustainable as peoples’ lives advance?
@@
Yet SAM (Chen et al., 2025b) remains primarily a conceptual proposition. As a speculative design framework, it offers a compelling vision but has not been operationalized through design practices, workshops, or any empirical inquiry. Our contribution is to build on this conceptual vision by making SAM actionable through modular co-design. Whereas SAM highlights the theoretical interdependence of sustainability, adaptability, and modularity, our work translates these principles into design practice. Additionally, while SAM positions sustainability, adaptability, and modularity as coequal and mutually reinforcing, our work deliberately centers on modularity as a participatory practice that can enact and extend SAM in lived contexts. To further extend SAM into real-world user-centered design, we explicitly incorporate personalization: it allows people to tailor technologies to their own needs and identities, making robots more meaningful in everyday life.
Specifically, we reframe modularity in our design framework not only as a purely technical property of reconfigurable robots (Yim et al., 2000; Seo et al., 2019; Zou et al., 2022; Deniša et al., 2023), but as a participatory practice in which users actively assemble, disassemble, and reconfigure interchangeable modules. This approach empowers users to personalize their robots, supports flexible adaptation over time, and embeds repair, reuse, and upgrading into everyday interaction (Maestri and Wakkary, 2011). Thus, modular co-design not only aligns with SAM but extends it, operationalizing the framework by granting users agency and rooting sustainability, adaptability, and modularity in lived practices of co-creation. In this paper, we ask: How can modular co-design support personalization, adaptability, and sustainability across the human lifespan in future robot design?
To explore this question, we conducted two co-design workshops with 15 younger adults (ages 21–32) and 8 older adults (65+). Participants were asked to imagine how a single modular robot could accompany them from childhood through adulthood and into older adulthood, reconfiguring through interchangeable modules (heads, arms, torsos, and bases) to meet changing needs. Our findings highlight three mutually reinforcing qualities of modular design. Personalization enabled participants to configure robots that reflected their unique values and circumstances, even within the same life stage. Adaptability allowed these robots to flexibly shift roles, both within a single stage of life (e.g., balancing work and care in adulthood) and across stages (e.g., from playful peer in childhood to supportive companion in older adulthood). Sustainability emerged at the intersection of these qualities: (a) materially, through repair, reuse, and upgrading of modules; and (b) emotionally, through continuity, attachment, and identity over time. These findings show how modularity, which enables user co-design robot practices, transforms robots from disposable gadgets into enduring companions.
@@
We advance conceptual understanding by positioning personalization, adaptability, and sustainability as mutually reinforcing dimensions rather than separate concerns.
@@ 2. Related Work
In this section, we review two strands of prior work that inform our study: (1) research on sustainability in Human–Robot Interaction and the emerging role of modular robotics (see subsection 2.1), and (2) research on robots across the human lifespan (see subsection 2.2). These perspectives highlight both the ecological and experiential dimensions of sustainability, and point to the need for user-centered approaches that envision robots adapting with people over time.
@@ 2.1. Sustainable HRI and Modular Robots
Concerns about sustainability have long shaped Human–Computer Interaction (HCI), where frequent cycles of consumption and technological obsolescence generate significant environmental impacts (Cooper, 2012; DiSalvo et al., 2010; Hansson et al., 2021; Remy and Huang, 1408; Yu et al., 2024). In 2007, Blevis introduced the concept of Sustainable Interaction Design (SID), also called Sustainable HCI (SHCI) (Hansson et al., 2021), calling sustainability to be a central concern of design (Blevis, 2007). Blevis’s framework (Blevis, 2007) emphasized that interaction design choices have material consequences, from rare metal mining and energy consumption to the accumulation of toxic electronic waste.
@@
Yet, most current robots are designed as rigid, closed systems (Šabanović, 2010). They cannot be easily upgraded, adapted, or repaired, leading to cycles of replacement and waste. Modular robotics recently offers a promising alternative (Alattas et al., 2019). By enabling individual components to be swapped, recombined, or repaired, modular systems can extend the useful life of a robot while supporting personalization and iterative adaptation (Alattas et al., 2019; Brunete et al., 2017). Technical research has already shown capacities such as self-assembly, self-reconfiguration, self-repair, and self-reproduction (Alattas et al., 2019). Classic examples include PolyBot, which could reconfigure itself to adapt to different terrains (Yim et al., 2000); Roombots, which transform into furniture for reuse in everyday contexts (Sproewitz et al., 2009; Spröwitz et al., 2010); and EMeRGE, an open modular platform that supports continuous experimentation and incremental innovation (Moreno and Faiña, 2021). These systems illustrate how modularity underpins resilience and long-term adaptability in robotics. However, sustainability in robotics should not only be understood in terms of material efficiency or waste reduction. A truly sustainable approach also requires attending to how robots remain meaningful, adaptable, and usable for people over time (Gajšek et al., 2020). Despite technical advances, modular robotics research has largely focused on engineering feasibility, with limited attention to how modularity aligns with human needs, values, and practices across the lifespan.
@@
To address this gap, we draw on the robot design theory of Sustainability, Adaptability, and Modularity (SAM) (Chen et al., 2025b). SAM (Chen et al., 2025b) conceptualizes Sustainability, Adaptability, and Modularity as three mutually reinforcing pillars of future robot design. It positions sustainability as the overarching aim, adaptability as the robot’s capacity to evolve with people, and modularity as the mechanism enabling repair and upgrading. However, SAM remains primarily a speculative design framework: it outlines a conceptual constellation but does not provide methods for engaging users or empirically grounding these principles.
@@
By contrast, our work proposes a distinct framework of Modular Co-Design for Personalization, Adaptability, and Sustainability. This framework differs from SAM in two key ways. First, it foregrounds modularity as participatory practice, treating users’ active assembly, disassembly, and reconfiguration as the mechanism through which adaptability and sustainability can be realized. Second, it foregrounds personalization as a user-centered dimension, underscoring the ways in which people modify and adjust robots to fit their changing circumstances across life stages that SAM only hints at rather than fully develops.
@@ 2.2. Robots Across the Lifespan
HRI studies showed that for robots to be genuinely effective in people’s lives, they should do more than complete short-term tasks (Dautenhahn, 2007; Kidd and Breazeal, 2008). They need to adapt to a wide range of situations and provide personalized support that evolves with users over time (Umbrico et al., 2020). Yet, as Irfan et al. notes, most HRI research has focused on brief interactions, even though many real-world applications, such as caregiving, education, and companionship, require long-term relationship building between humans and robots (Irfan et al., 2023, 2019). Existing studies often isolate user groups into narrow categories, such as children or older adults, examining them separately (Ahmad et al., 2017; Coninx et al., 2016; Yamazaki et al., 2023; Ostrowski et al., 2022). While these segmented approaches offer valuable insights into age-specific needs, they tend to overlook how expectations change as people move through different life stages. They also miss the deeper emotional attachments that can develop when robots remain present across longer periods of time. This gap underscores the importance of design approaches that can support sustainable, adaptable, and emotionally resonant engagement throughout the human lifespan.
@@
Understanding how robots can adapt fluidly over the lifespan would not only guide the design of long-term interaction strategies, but also highlight the potential for sustained human–robot attachment (Kanda et al., 2007; Leite et al., 2013). Such continuity can reinforce sustainability by reducing abandonment, extending product lifespans, and minimizing the need for complete system replacements (Kamino, 2023). In this sense, modular robots hold the potential to achieve what some scholars describe as “heirloom status” (Türe and Ger, 2016), wherein a technology accumulates emotional and symbolic value across generations. By retaining familiar modules and storing shared histories, robots may undergo a process of “ensoulment” (Blevis1 and Stolterman1, 2007; Jung et al., 2011; Nelson and Stolterman, 2014; Odom et al., 2009), evolving into more than digital devices: they become companions imbued with personal and familial meaning. Much like a grandmother’s wedding ring, a modular robot could one day be cherished and even passed down, carrying both material continuity and emotional resonance.
Our study builds directly on this perspective. By engaging participants in imagining a single modular robot that could accompany them from childhood to adulthood and into older age, we contribute empirical insights into what “lifespan sustainability” might look like in practice. Specifically, we show how modularity can serve as the mechanism that enables adaptation, how adaptation fosters long-term emotional attachment, and how this combination supports sustainability in both ecological and experiential terms. In doing so, our work extends the discussion of robot lifespans from abstract aspiration to concrete user-centered visions, highlighting design opportunities for creating robots that endure, not just as tools, but as lifelong companions.
@@ 3. Methods: Co-Design Workshop about Modular Robots Across Lifespan
Co-design means designing something together with the people who might one day use it. Instead of asking people only to test a finished product, researchers invite them to take part from the very beginning: sharing ideas, imagining possible futures, and helping shape how a system should look and work. This approach is especially valuable in HRI research (Burlando et al., 2021; Gebell´ı et al., 2024), where robots are often expected to support people over long periods of time and in personal settings such as homes (Hofstede et al., 2024; Ostrowski et al., 2020; Chen et al., 2025a). By making future users active creative partners, co-design helps ensure that robot designs fit people’s real needs and experiences (Hsu et al., 2024; Garg and Cui, 2022).
In our study, we used co-design to explore how modular robots could grow and change alongside people at different life stages—childhood, adulthood, and older adulthood. To do so, we conducted two in-person workshops in the United States: one with 15 younger adults (ages 21–32) and another with 8 older adults (ages 66–81). This dual-group design allowed us to compare forward-looking imaginations from younger adults with experience-informed reflections from older adults, thereby combining speculative and experiential perspectives on lifespan-oriented modular robotics.
To guide both workshops, we adopted the robot modules and functions proposed by Chen et al. in the 2025 ACM Conference on Creativity and Cognition (see Figure 1, Figure 2, Figure 3, Figure 4). This framework is based on interchangeable building blocks, such as heads (H), arms (A), upper bodies (UB), and lower bodies (LB), that can be swapped or combined (see Figure 6 about specific module robot design and functions). Such modularity allows for personalization, repair, and reuse, while also giving participants freedom to creatively configure robots that match their changing needs across a lifetime.
@@ 3.1. Participants
black
Participant
Gender
Age
Education
Robot Experience
P1
M
24
Masters
Have limited experience
P2
M
21
High School
Have used in everyday life
P3
F
27
Masters
Have limited experience
P4
F
27
Masters
Have limited experience
P5
F
32
Masters
Have limited experience
P6
F
25
Masters
Have never used
P7
F
28
Ph.D.
Have used in everyday life
P8
M
24
Bachelors
Have limited experience
P9
F
24
Masters
Have limited experience
P10
F
29
Ph.D.
Have used in everyday life
P11
F
27
Masters
Have limited experience
P12
M
26
Masters
Have limited experience
P13
F
26
Masters
Have limited experience
P14
F
26
Masters
Have limited experience
P15
F
30
Masters
Have never used
P16
F
73
Masters
Have used in everyday life
P17
M
73
Bachelors
Have never used
P18
M
71
Ph.D.
Have limited experience
P19
F
68
Bachelors
Have limited experience
P20
F
67
Masters
Have never used
P21
F
75
Bachelors
Have used in everyday life
P22
F
66
Masters
Have limited experience
P23
F
81
Ph.D.
Have never used
@@
The first workshop involved 15 younger adults (P1 to P15), aged 21 to 32 (M = 26.1, SD = 3.3), recruited through a local university. Participants came from varied academic and professional backgrounds, including Human–Computer Interaction, Computer Science, Cyber-security, and Business Management. They represented diverse educational backgrounds, from high school to Ph.D. Younger adults, who have not yet experienced later life stages, could only imagine future needs through aspiration and projection. Their perspectives were therefore largely forward-looking, speculative, and shaped by anticipated trajectories rather than lived realities.
To complement this group, we conducted a second workshop with 8 older adults (P16 to P23), aged 66 to 81 (M = 72.9, SD = 5.2), recruited through senior centers and community networks. While they also had no direct experience with modular robot design, their ideas were grounded in everyday realities of aging, caregiving, and long-term domestic support. In contrast to the younger group’s aspirational orientation, older adults provided reflections informed by firsthand experiences of later life challenges and support needs.
By comparing these two groups under a unified design process, our study combines forward-looking imagination with experience-based reflection. This dual perspective enables us to capture both the speculative possibilities and the grounded constraints of lifespan-oriented modular robots. Demographic details and prior robot experiences of all 23 participants are summarized in Table 1.
@@
| H3 | The triangular shape helps to optimizes the movement on unstable terrain. |
@@
| H7 | Scans the person’s head shape and use virtual reality to display different hairstyle options. |
| H8 | Grows and sustain real plants, provides optimal care. |
@@
| A6 | Suitable for hospitals or home care, it can hold items such as IV bags or deliver medication. |

### v2

@@ 1. Introduction
Robots are increasingly envisioned not as short-term tools but as long-term companions that support people throughout their lives (Leite et al., 2013; Kaplan, 2005; Chen et al., 2025a). From childhood education to workplace productivity and eldercare (Chen et al., 2025a; Bemelmans et al., 2012; Kory-Westlund and Breazeal, 2019; Westlund et al., 2018; Yi et al., 2025), socially assistive robots hold the promise of providing functional help (Deutsch et al., 2019; Klamer and Allouch, 2010), emotional connection (Šabanović et al., 2013a), and continuity of care across the human lifespan (Irfan et al., 2023; Kidd and Breazeal, 2008). Realizing this promise, however, requires moving beyond one-size-fits-all systems. Because people’s needs, values, and contexts shift over time (Carstensen et al., 1999; Fung et al., 2001), robots should remain not only useful at one life stage but also adaptable over decades (De Graaf et al., 2015). This raises a central challenge for Human–Robot Interaction (HRI): how can robots remain relevant, meaningful, and sustainable as peoples’ lives advance?
@@
While SAM (Chen et al., 2025b) remains a speculative, system-level design theory focused on ideal robot properties, we build on this work by naming a second complementary framework. Herein, we introduce the Personalization, Adaptability, Sustainability (PAS) framework as a complementary, human-centered model derived from empirical co-design, specifically capturing the lived experience of lifespan HRI. PAS reframes its own three dimensions from within participants’ imagined, ongoing relationship with a single modular robot: personalization captures how people use modules to inscribe their own roles, preferences, and values into the robot; adaptability describes how they reconfigure its form and functions as life events and responsibilities shift; and sustainability names the material and emotional continuity that emerges when modules are repaired, reused, and carried forward over time. In this view, modularity is no longer a separate pillar but the socio-technical practice of assembling, swapping (Seo et al., 2019; Zou et al., 2022; Deniša et al., 2023), and recombining parts through which sustainability, adaptability, and personalization are actually enacted in people’s lives.
In this paper, we ask: How can modular co-design support personalization, adaptability, and sustainability across the human lifespan in future robot design? To explore this question, we conducted two co-design workshops with 15 younger adults (ages 21–32) and 8 older adults (65+). Participants were asked to imagine how a single modular robot could accompany them from childhood through adulthood and into older adulthood, reconfiguring through interchangeable modules (heads, arms, torsos, and bases) to meet changing needs. Our findings highlight three mutually reinforcing qualities of modular design. Personalization enabled participants to configure robots that reflected their unique values and circumstances, even within the same life stage. Adaptability allowed these robots to flexibly shift roles, both within a single stage of life (e.g., balancing work and care in adulthood) and across stages (e.g., from playful peer in childhood to supportive companion in older adulthood). Sustainability emerged at the intersection of these qualities: (a) materially, through repair, reuse, and upgrading of modules; and (b) emotionally, through continuity, attachment, and identity over time. These findings show how modularity, which enables user co-design robot practices, transforms robots from disposable gadgets into enduring companions.
@@
We advance a conceptual understanding of HRI durability by introducing an additional human-centered framework, namely Personalization, Adaptability, and Sustainability (PAS) as a complementary framework to SAM. PAS adds focus on the experiential dimensions of long-term human-robot relationships.
@@ 2. Related Work
In this section, we review two strands of prior work that inform our study: (1) research on sustainability in Human–Robot Interaction and the emerging role of modular robotics (see subsection 2.1), and (2) research on robots across the human lifespan (see subsection 2.2). These perspectives highlight both the ecological and experiential dimensions of sustainability, and point to the need for human-centered approaches that envision robots adapting with people over time.
@@ 2.1. Sustainable HRI and Modular Robots
Concerns about sustainability have long shaped Human–Computer Interaction (HCI), where frequent cycles of consumption and technological obsolescence generate significant environmental impacts (Cooper, 2012; DiSalvo et al., 2010; Hansson et al., 2021; Remy and Huang, 2014; Yu et al., 2024). In 2007, Blevis introduced the concept of Sustainable Interaction Design (SID), also called Sustainable HCI (SHCI) (Hansson et al., 2021), calling sustainability to be a central concern of design (Blevis, 2007). Blevis’s framework (Blevis, 2007) emphasized that interaction design choices have material consequences, from rare metal mining and energy consumption to the accumulation of toxic electronic waste.
@@
Yet, most current robots are designed as rigid, closed systems (Šabanović, 2010; Koike et al., 2024; Tang et al., 2025). They cannot be easily upgraded, adapted, or repaired, leading to cycles of replacement and waste. Modular robotics recently offers a promising alternative (Alattas et al., 2019). By enabling individual components to be swapped, recombined, or repaired, modular systems can extend the useful life of a robot while supporting personalization and iterative adaptation (Alattas et al., 2019; Brunete et al., 2017). Technical research has already shown capacities such as self-assembly, self-reconfiguration, self-repair, and self-reproduction (Alattas et al., 2019). Classic examples include PolyBot, which could reconfigure itself to adapt to different terrains (Yim et al., 2000); Roombots, which transform into furniture for reuse in everyday contexts (Sproewitz et al., 2009; Spröwitz et al., 2010); and EMeRGE, an open modular platform that supports continuous experimentation and incremental innovation (Moreno and Faiña, 2021). These systems illustrate how modularity underpins resilience and long-term adaptability in robotics. However, sustainability in robotics should not only be understood in terms of material efficiency or waste reduction. A truly sustainable approach also requires attending to how robots remain meaningful, adaptable, and usable for people over time (Gajšek et al., 2020). Despite technical advances, modular robotics research has largely focused on engineering feasibility, with limited attention to how modularity aligns with human needs, values, and practices across the lifespan.
@@
To address this gap, we draw on the robot design theory of Sustainability, Adaptability, and Modularity (SAM) (Chen et al., 2025b). SAM (Chen et al., 2025b) conceptualizes Sustainability, Adaptability, and Modularity as three mutually reinforcing pillars of future robot design. Beyond this conceptual framing, our module set also builds on the collection of interchangeable robot parts and a configurable “robot ecology” of heads, arms, bodies, and bases (Chen et al., 2025b). However, SAM remains primarily a speculative design framework: it outlines a conceptual constellation but does not provide methods for engaging users or empirically grounding these principles.
@@
By contrast, our work proposes a distinct, human-centered framework of Modular Co-Design for Personalization, Adaptability, and Sustainability (PAS). This framework differs from SAM in two key ways. First, it foregrounds modularity as participatory practice, treating users’ active assembly, disassembly, and reconfiguration as the mechanism through which adaptability and sustainability can be realized. Second, it foregrounds personalization as a human-centered dimension, underscoring the ways in which people modify and adjust robots to fit their changing circumstances across life stages that SAM only hints at rather than fully develops. Personalization, in this sense, is not an additional feature but a theoretical correction: without attending to how robots are tuned to particular people and contexts, the promise of sustainability and adaptability risks remaining abstract and normative.
@@ 2.2. Robots Across the Lifespan
HRI studies showed that for robots to be genuinely effective in people’s lives, they should do more than complete short-term tasks (Dautenhahn, 2007; Kidd and Breazeal, 2008; Tang et al., 2025). They need to adapt to a wide range of situations and provide personalized support that evolves with users over time (Umbrico et al., 2020; Tang et al., 2025; Han et al., 2024, 2023). Yet, as Irfan et al. notes, most HRI research has focused on brief interactions, even though many real-world applications, such as caregiving, education, and companionship, require long-term relationship building between humans and robots (Irfan et al., 2023, 2019). Existing studies often isolate user groups into narrow categories, such as children or older adults, examining them separately (Ahmad et al., 2017; Coninx et al., 2016; Yamazaki et al., 2023; Ostrowski et al., 2022). While these segmented approaches offer valuable insights into age-specific needs, they tend to overlook how expectations change as people move through different life stages. They also miss the deeper emotional attachments that can develop when robots remain present across longer periods of time. Recent work has further explored how robots might embody long-term temporal change through material processes such as growth, aging, and decay, offering alternative interaction trajectories that unfold over days or months rather than minutes (Hu et al., 2024). This gap underscores the importance of design approaches that can support sustainable, adaptable, and emotionally resonant engagement throughout the human lifespan.
@@
Understanding how robots can adapt fluidly over the lifespan would not only guide the design of long-term interaction strategies, but also highlight the potential for sustained human–robot attachment (Kanda et al., 2007; Leite et al., 2013). Such continuity can reinforce sustainability by reducing abandonment, extending product lifespans, and minimizing the need for complete system replacements (Kamino, 2023). In this sense, modular robots hold the potential to achieve what some scholars describe as “heirloom status” (Blevis, 2007; Jung et al., 2011), wherein a technology accumulates emotional and symbolic value across generations. By retaining familiar modules and storing shared histories, robots may undergo a process of “ensoulment” (Blevis1 and Stolterman1, 2007; Nelson and Stolterman, 2014; Odom et al., 2009), evolving into more than digital devices: they become companions imbued with personal and familial meaning. Much like a grandmother’s wedding ring, a modular robot could one day be cherished and even passed down, carrying both material continuity and emotional resonance.
Our study builds directly on this perspective. By engaging participants in imagining a single modular robot that could accompany them from childhood to adulthood and into older age, we contribute empirical insights into what “lifespan sustainability” might look like in practice. Specifically, we show how modularity can serve as the mechanism that enables adaptation, how adaptation fosters long-term emotional attachment, and how this combination supports sustainability in both ecological and experiential terms. In doing so, our work extends the discussion of robot lifespans from abstract aspiration to concrete human-centered visions, highlighting design opportunities for creating robots that endure, not just as tools, but as lifelong companions.
@@ 3. Methods: Co-Design Workshop about Modular Robots Across Lifespan
Co-design means designing something together with the people who might one day use it. Instead of asking people only to test a finished product, researchers invite them to take part from the very beginning: sharing ideas, imagining possible futures, and helping shape how a system should look and work. This approach is especially valuable in HRI research (Burlando et al., 2021; Gebellí et al., 2024), wherein robots are often expected to support people over long periods of time and in personal settings such as homes (Hofstede et al., 2024; Ostrowski et al., 2020; Chen et al., 2025a). By making future users active creative partners, co-design helps ensure that robot designs fit people’s real needs and experiences (Hsu et al., 2024; Garg and Cui, 2022). Our study contributes a specific variant of this approach: we use a modular set of physical robot parts as a temporal design probe, inviting participants to repeatedly reconfigure the same robot across imagined life stages. In doing so, modularity itself becomes a research medium for observing how preferences, expectations, and repair practices unfold over a hypothesized lifespan.
In our study, we used co-design to explore how modular robots could grow and change alongside people at different life stages: childhood, adulthood, and older adulthood. To do so, we conducted two in-person workshops in the United States: one with 15 younger adults (ages 21–32) and another with 8 older adults (ages 66–81). This dual-group design allowed us to compare forward-looking imaginations from younger adults with experience-informed reflections from older adults, thereby combining speculative and experiential perspectives on lifespan-oriented modular robotics.
To guide both workshops, we adopted the robot modules and functions proposed by Chen et al. in the 2025 ACM Conference on Creativity and Cognition (see Figure 1, Figure 2, Figure 3, Figure 4). This framework is based on interchangeable building modules, such as heads (H), arms (A), upper bodies (UB), and lower bodies (LB), that can be swapped or combined. We initially kept this structure and four body-like grouping as a starting assumption rather than an end goal—a familiar construction that we expected participants to take up, question, or break. We chose these modules for two reasons. First, they offered a ready-made, functionally diverse toolkit that allowed participants to focus on imagining how roles and relationships change across the lifespan rather than on inventing entirely new mechanical parts from scratch. Second, the familiar “body-like” groupings provided an accessible starting point for participants with no robotics background, while still allowing for playful recombination and distortion.
Importantly, we treated this anthropomorphic scaffold as a starting assumption rather than an end goal. Participants were explicitly encouraged to rearrange, duplicate, or omit parts and to experiment with non-humanoid configurations. As we discuss in our findings, many initially assembled recognizably humanoid robots but then stretched or violated the template—stacking multiple heads, discarding legs, or creating distributed, multi-body systems. This tension between starting from a human-like structure and then breaking it is central to our analysis: it shows how modularity can both reproduce default expectations about robot form and provide a resource for challenging them.
@@ 3.1. Participants
| Participant | Gender | Age | Education | Robot Experience |
| P1 | M | 24 | Masters | Have limited experience |
| P2 | M | 21 | High School | Have used in everyday life |
| P3 | F | 27 | Masters | Have limited experience |
| P4 | F | 27 | Masters | Have limited experience |
| P5 | F | 32 | Masters | Have limited experience |
| P6 | F | 25 | Masters | Have never used |
| P7 | F | 28 | Ph.D. | Have used in everyday life |
| P8 | M | 24 | Bachelors | Have limited experience |
| P9 | F | 24 | Masters | Have limited experience |
| P10 | F | 29 | Ph.D. | Have used in everyday life |
| P11 | F | 27 | Masters | Have limited experience |
| P12 | M | 26 | Masters | Have limited experience |
| P13 | F | 26 | Masters | Have limited experience |
| P14 | F | 26 | Masters | Have limited experience |
| P15 | F | 30 | Masters | Have never used |
| P16 | F | 73 | Masters | Have used in everyday life |
| P17 | M | 73 | Bachelors | Have never used |
| P18 | M | 71 | Ph.D. | Have limited experience |
| P19 | F | 68 | Bachelors | Have limited experience |
| P20 | F | 67 | Masters | Have never used |
| P21 | F | 75 | Bachelors | Have used in everyday life |
| P22 | F | 66 | Masters | Have limited experience |
| P23 | F | 81 | Ph.D. | Have never used |
@@
The first workshop involved 15 younger adults (P1 to P15), aged 21 to 32 (M = 26.1, SD = 3.3), recruited through a local university. Participants came from varied academic and professional backgrounds, including Human–Computer Interaction, Computer Science, Cyber-security, and Business Management. They represented diverse educational backgrounds, from high school to Ph.D. Younger adults, who have not yet experienced later life stages, could only imagine future needs through aspiration and projection. Their perspectives mainly combined forward-looking imagination with memories of their own childhood, and occasionally with experiences of being around younger relatives or children.
To complement this group, we conducted a second workshop with 8 older adults (P16 to P23), aged 66 to 81 (M = 72.9, SD = 5.2), recruited through senior centers and community networks. While they also had no direct experience with modular robot design, their ideas were grounded in everyday realities of aging, caregiving, and long-term domestic support. In contrast to the younger group’s aspirational orientation, older adults provided reflections informed by firsthand experiences of later life challenges and support needs, and drew on a mix of autobiographical memories and occasionally their experiences as parents or grandparents when reasoning about childhood.
By comparing these two groups under a unified design process, our study combines forward-looking imagination with experience-based reflection. This dual perspective enables us to capture both the speculative possibilities and the grounded constraints of lifespan-oriented modular robots. Demographic details and prior robot experiences of all 23 participants are summarized in Table 1. Beyond these self-reported categories, participants who have experience with robots described prior experiences primarily with domestic service robots (e.g., robotic vacuum cleaners), educational robots used in classes or labs, and voice-based smart assistants. Only a small subset mentioned prior exposure to explicitly “companion-like” robots (e.g., pet-like or socially assistive systems), and those encounters were typically short and institutional (e.g., museums, care facilities) rather than long-term in the home.
@@
| H3 | The triangular shape helps to optimize the movement on unstable terrain. |
@@
| H7 | Scans the person’s head shape and uses virtual reality to display different hairstyle options. |
| H8 | Grows and sustains real plants, provides optimal care. |
@@
| A6 | Suitable for hospitals or home care, which can hold items such as IV bags or deliver medication. |